OpenStudy (anonymous):

I need help where to start on this question i think what confuses me from the get go is the exponents, the question ask to express in trignometric form. Z=(sqrt 3 - i)^2/(sqrt 2 - i sqrt 2)^6

OpenStudy (anonymous):

Let z = a + bi. Then r = sqrt(a^2+b^2) = sqrt(1^2+(sqrt3)^2) = 2 and phi = arctan(b/a) = arctan(sqrt3) = 60° = pi/3 So you have polar form: z = r*e^(i*phi) = 2*e^(i*pi/3) and trigonometric form: z = r*(cos(phi) + i*sin(phi)) = 2*(cos(pi/3) + i*sin(pi/3))

OpenStudy (anonymous):

that would be for the numerator part where you changed it into trig form, but afterwards would i apply z1//z2 quotient formula

OpenStudy (anonymous):

http://gamutech.com